OpenStudy (anonymous):

Let V be a vector space and ℓ : V → R be a linear map. If z ∈ V is not in the nullspace of ℓ, show that every x ∈ V can be decomposed uniquely as x = v + cz, where v is in the nullspace of ℓ and c is a scalar.

OpenStudy (anonymous):

its some crazy stuff though.

OpenStudy (dan815):

is it interesting stuff

OpenStudy (anonymous):

yea i like it more than calculus

OpenStudy (dan815):

ohh really!! i want to learn then

OpenStudy (dan815):

why do u like it more than calculus?

OpenStudy (anonymous):

there exists weird things in algebra...but calculus from my experience is mostly alot of methodology...more ambiguity in linear algebra.

OpenStudy (anonymous):

like why is there called an empty set which contains nothing..

terenzreignz (terenzreignz):

I don't see questions like these often :D

OpenStudy (anonymous):

does that excite you?

OpenStudy (dan815):

it excites me

terenzreignz (terenzreignz):

a little :)

terenzreignz (terenzreignz):

Well, @dan815 after you :)

OpenStudy (dan815):

i am excited for different reasons than u

OpenStudy (dan815):

iam excited because i dont understand these words haHahaha

OpenStudy (anonymous):

hahaha. dan atm has more of an idea how cows eat grass than this question right now

terenzreignz (terenzreignz):

Oh, make no mistake, I DO NOT know the answer here... (yet) but I'm working on it :)

terenzreignz (terenzreignz):

I suppose we start with the essentials. Could I use f instead of ℓ ? It's easier to type it that way :)

OpenStudy (anonymous):

yea k.

terenzreignz (terenzreignz):

okay, the usuals... Let V be a vector space, and f: V -> R be a linear map. Let z be an element of v such that f(v) is not equal to 0

terenzreignz (terenzreignz):

And now, let x be an element of V.

OpenStudy (anonymous):

yepyepyep

OpenStudy (anonymous):

That's a HARD one!

OpenStudy (anonymous):

AND UR HERE TO HELP :) haha

OpenStudy (anonymous):

Haha Yes , Yes I am :D

OpenStudy (anonymous):

goodie !

terenzreignz (terenzreignz):

Thankfully, the co-domain is just the set of reals. I think we can do this now... :D

OpenStudy (anonymous):

Im still tryna figure it out here.. HM.......

terenzreignz (terenzreignz):

Why don't we get rid of the annoying trivial case of when x itself is in the kernel (nullspace) of f. Then it follows trivially, just set v = x and c = 0

OpenStudy (anonymous):

alright...mmm. i've got till 11:30pm so im 1hour

OpenStudy (anonymous):

Great!

terenzreignz (terenzreignz):

relax :P Now what if x is not in the (bloody) nullspace ? Well then, we can set \[\Large c = \frac{f(x)}{f(z)}\]

OpenStudy (anonymous):

haha

terenzreignz (terenzreignz):

Which exists, because f(z) is not in the nullspace (f(z) is not zero)

terenzreignz (terenzreignz):

\[\Large\qquad \ \ f(x) = cf(z)\]\[\Large \implies f(x)= f(cz)\]

OpenStudy (anonymous):

oH I WAS AT THE WROND dESTINATION^

terenzreignz (terenzreignz):

(remembering of course that f is a linear transformation)

terenzreignz (terenzreignz):

Okay, that said, we set \[\Large v = x - cz\]

terenzreignz (terenzreignz):

Let's TRANSFORM both sides... \[\Large f(v) = f(x-cz) = f(x)-f(cz) = 0\]

OpenStudy (anonymous):

i'm with ya

OpenStudy (anonymous):

me too

terenzreignz (terenzreignz):

And you have \[\Large f(v) = 0\] implying, of course, that v is in the nullspace of f.

OpenStudy (anonymous):

haha zofia. let terenz do all the talking :)

OpenStudy (anonymous):

okk

terenzreignz (terenzreignz):

-.- Anyway, \[\Large v = x - cz\] Then \[\Large x = v + cz\] Successfully decomposing x as desired. Now, maybe you can show uniqueness?

OpenStudy (anonymous):

oK cHRIS, sowwie!

OpenStudy (anonymous):

nono. i didn't mean it like that! like ur learning too:)

OpenStudy (anonymous):

so we are both students:) haha yea ok.

OpenStudy (anonymous):

riiigghht! :D

terenzreignz (terenzreignz):

uhh, guys? focus? :D Showing uniqueness... any idea guys? :)

OpenStudy (anonymous):

um. let me think

terenzreignz (terenzreignz):

Here, let me start you out :) Suppose we have two such decompositions of x... \[\Large x = v_1 + c_1z\]\[\Large x = v_2 + c_2z\] You have to show that in fact, \[\Large v_1 =v_2\]and\[\Large c_1=c_2\]

OpenStudy (anonymous):

equality of both sides?

terenzreignz (terenzreignz):

Equality, specifically, of v1 and v2 as well as c1 and c2.

OpenStudy (anonymous):

yea

terenzreignz (terenzreignz):

Where both \(v_1\) and \(v_2\) are in the nullspace of f and \(c_1\) and \(c_2\) are scalars

terenzreignz (terenzreignz):

Stuck? :)

OpenStudy (anonymous):

u read mi mind

terenzreignz (terenzreignz):

You could start with this... remembering that since both expressions are equal to x, then you have \[\Large v_1 + c_1z = v_2 + c_2z\] Could you work from there?

OpenStudy (anonymous):

that wat i meant by equality of both sides...

OpenStudy (anonymous):

so c1=c2 from z:

OpenStudy (anonymous):

I get it !

terenzreignz (terenzreignz):

I'm afraid it's not that simple, as you haven't established equality between c1 and c2 yet :P

OpenStudy (anonymous):

o poo

OpenStudy (anonymous):

i forgot abotu that

terenzreignz (terenzreignz):

@Zofia1 yes? :)

OpenStudy (anonymous):

Yes.

terenzreignz (terenzreignz):

I meant, you know how to do it?

OpenStudy (anonymous):

Im doing on paper and here wait.

terenzreignz (terenzreignz):

No rush.

OpenStudy (anonymous):

you group c1z an c2z on one side?

terenzreignz (terenzreignz):

Yup :) \[\Large v_1 - v_2 = c_2z - c_1z\]

OpenStudy (anonymous):

and factor?

terenzreignz (terenzreignz):

Very intuitive :) \[\Large v_1 - v_2 = (c_2-c_1)z\]

OpenStudy (anonymous):

then we estalibished f(v)=0?

terenzreignz (terenzreignz):

hmm... almost, not quite :)

terenzreignz (terenzreignz):

Remember, we have a linear transformation f, where both v1 and v2 are elements of its nullspace and z specifically isn't...

OpenStudy (anonymous):

does this mean v1=0 and v2=0? because they are in the nullspace

terenzreignz (terenzreignz):

Don't be in such a rush :D Let's apply the linear transformation f on both sides, shall we? \[\Large f(v_1 - v_2 )= f[(c_2-c_1)z]\]

terenzreignz (terenzreignz):

Now, on the left, f being a linear transformation, we can have... \[\Large f(v_1) - f(v_2 )= f[(c_2-c_1)z]\] With me so far?

OpenStudy (anonymous):

yea. i get that.

terenzreignz (terenzreignz):

On the right, since c2 - c1 is just a scalar, likewise, we can simplify into... \[\Large f(v_1) - f(v_2 )= (c_2-c_1)f(z)\]

OpenStudy (anonymous):

cant we do that to the right hand side?

OpenStudy (anonymous):

o isee

terenzreignz (terenzreignz):

c2 - c1 is a scalar...

OpenStudy (anonymous):

OH!

terenzreignz (terenzreignz):

Okay, ready to proceed? Or can you demonstrate it from here?

OpenStudy (anonymous):

I MEAN OH!

terenzreignz (terenzreignz):

hmm?

OpenStudy (anonymous):

left hand side is 0 because we said f(v)=0 right?

terenzreignz (terenzreignz):

Left hand side is zero because BOTH \(\large f(v_1)\) and \(\large f(v_2)\) are zero, as both v1 and v2 are in the nullspace

OpenStudy (anonymous):

now f(z) is not in the nullspace so that means c2-c1 must only equal zero.

OpenStudy (anonymous):

for this to occur c2=c1

terenzreignz (terenzreignz):

Yes continue... (you haven't shown equality between v1 and v2 quite yet...)

OpenStudy (anonymous):

ok. so we showed equality with the scalars.

terenzreignz (terenzreignz):

Right...

OpenStudy (anonymous):

we can still use this yea? f(v1)−f(v2)=(c2−c1)f(z)

OpenStudy (anonymous):

f(v1−v2)=f[(c2−c1)z]

OpenStudy (anonymous):

that one